# Towards Heterogeneous Continual Graph Learning via Meta-knowledge Distillation

### v1

@@
Abstract
Machine learning on heterogeneous graphs has experienced rapid advancement in recent years, driven by the inherently heterogeneous nature of real-world data. However, existing studies typically assume the graphs to be static, while real-world graphs are continuously expanding. This dynamic nature requires models to adapt to new data while preserving existing knowledge. To this end, this work addresses the challenge of continual learning on heterogeneous graphs by introducing the Meta-learning based Knowledge Distillation framework (MKD), designed to mitigate catastrophic forgetting in evolving heterogeneous graph structures. MKD combines rapid task adaptation through meta-learning on limited samples with knowledge distillation to achieve an optimal balance between incorporating new information and maintaining existing knowledge. To improve the efficiency and effectiveness of sample selection, MKD incorporates a novel sampling strategy that selects a small number of target-type nodes based on node diversity and maintains fixed-size buffers for other types. The strategy retrieves first-order neighbors along metapaths and selects important neighbors based on their structural relevance, enabling the sampled subgraphs to retain key topological and semantic information. In addition, MKD introduces a semantic-level distillation module that aligns the attention distributions over different metapaths between teacher and student models, encouraging semantic consistency beyond the logit level. Comprehensive evaluations across three benchmark datasets validate MKD’s effectiveness in handling continual learning scenarios on expanding heterogeneous graphs.
1 Introduction
@@
In recent years, heterogeneous graphs have emerged as a powerful structure for modeling complex real-world systems across diverse domains, including biological systems, social networks, and recommendation systems shi2016heterogeneous . Unlike homogeneous graphs with uniform node and edge types, heterogeneous graphs capture intricate relationships among multiple node and edge types, enabling more nuanced and comprehensive representations for the target systems sun2012mining . For example, in recommendation system research, the user-item networks contain user nodes, item nodes, and the purchase relationship. In a cellular system, the signalling network encodes the complex interaction among different types of entities including small molecules, proteins, and genes zitnik2018modeling . To extract meaningful insights from heterogeneous graphs, different Heterogeneous Graph Neural Networks (HGNNs) have been developed. These approaches mainly fall into two categories: metapath-based HGNNs fu2020magnn ; zhang2019heterogeneous ; wang2019heterogeneous ; schlichtkrull2018modeling ; dong2017metapath2vec ; sun2011pathsim ; yun2019graph ; yang2023simple ; zhu2024hagnn , which aggregate information over predefined meta-paths to mine the intricate relationship among the heterogeneous nodes and edges, and meta-path-free models hu2020heterogeneous ; lv2021we ; zhu2019relation ; hong2020attention ; zhang2022simple ; fu2023multiplex ; huo2025heterogeneous that automatically learn node interactions without preconfigured paths, offering enhanced adaptability but potentially sacrificing interpretability.
While HGNNs have demonstrated remarkable success across various tasks, they face fundamental challenges in dynamic environments in which the heterogeneous graphs are constantly evolving with significant data distribution shift. For example, in recommendation systems, the user-item graphs undergo continuous expansion through new products, users, and purchase patterns. In biological systems, the signalling networks also grow with the discovery of novel signalling pathways. In these scenarios, HGNNs are expected to continuously incorporate the incoming new knowledge while preserving the learnt patterns. However, naive incremental training approaches will inevitably trigger catastrophic forgetting over previously learnt knowledge, as optimization for new data distributions tends to overwrite previously encoded information in the model parameters. While this phenomenon has been extensively studied in continual learning research over independent data (e.g. images) Aljundi_2018_ECCV ; kirkpatrick2017overcoming ; li2017learning ; lopez2017gradient and homogeneous graphs zhang2022cglb ; liu2021overcoming ; zhou2021overcoming ; zhang2023continual ; unal2023meta ; su2023towards ; ren2023incremental ; niu2024replay ; li2024efficient ; mondal2024stochastic ; liu2024towards , addressing catastrophic forgetting in heterogeneous graph learning remains largely unexplored.
@@
To bridge this critical gap, we present a systematic investigation of the Heterogeneous Continual Graph Learning (HCGL) problem and introduce Meta-learning based Knowledge Distillation (MKD), a novel framework specifically designed for HCGL. The proposed method consists of three core components: a Gradient-based Meta-learning Module (G-MM) for rapid adaptation, an Efficient Heterogeneous Subgraph Sampling (E-HSS) strategy for experience replay that leverages both node diversity and structural information, and a Heterogeneity-aware Knowledge Distillation (HKD) module that aligns cross-task knowledge at both the logit and semantic levels. First, unlike homogeneous graphs, the diverse node and edge types in heterogeneous graphs introduce extra challenges to knowledge transfer across different tasks. Therefore, for effective model adaptation over the continuously evolving heterogeneous graph structures, we design a Gradient-based Meta-learning Module (G-MM). G-MM learns an optimized initialization of model parameters by leveraging gradient-based meta-learning, which allows fast model adaptation to new tasks using only a small number of examples. Through task-specific adjustments, G-MM also avoids the task-wise interference and ensures performance robustness on learned tasks. Second, to preserve essential knowledge from previous tasks while avoiding excessive memory overhead, we propose an efficient Heterogeneous Subgraph Sampling (E-HSS) mechanism that integrates node diversity and structural information. E-HSS selects representative and important nodes from the target type and preserves the original topological structure guided by meta-paths. This strategy minimizes data redundancy by assigning independent buffers to different node types and computing node importance based on relation types and degree metrics, thereby achieving structural balance and semantic coverage across heterogeneous nodes and enhancing the model’s long-term memory retention. Finally, to align knowledge across tasks and maintain structural consistency in heterogeneous settings, we introduce a Heterogeneity-aware Knowledge Distillation (HKD) module. Beyond knowledge retained via experience replay, this module performs both logit-level and semantic-level alignment between the current and previous task models. By distilling prediction distributions and meta-path-based attention representations, HKD guides the student model to retain predictive information while incorporating multi-level semantic information, significantly improving continual learning performance on heterogeneous graphs. Our contribution can be summarized as follows:
  • We systematically investigate the problem of Heterogeneous Continual Graph Learning (HCGL), which has been largely unexplored in previous research.
@@
  • We propose an efficient sampling strategy that preserves both node diversity and heterogeneous structural information for memory-efficient experience replay.
@@
  • We introduce a two-level heterogeneity-aware distillation module that aligns knowledge across tasks at both the logit and semantic levels.
  • Extensive experiments on multiple benchmark heterogeneous graph datasets demonstrate that MKD significantly outperforms existing continual graph learning methods in terms of accuracy, efficiency, and memory utilization.
2 Related Work
@@
Heterogeneous Graph Neural Networks. Heterogeneous graphs are characterized by their complex topological structures and rich semantic information, stemming from the presence of multiple node and edge types. Traditional Graph Neural Networks (GNNs) often struggle to handle such heterogeneity directly, leading to the development of specialized approaches known as Heterogeneous Graph Neural Networks (HGNNs). These models are designed to effectively capture the multi-typed relationships inherent in heterogeneous graphs. Existing HGNN methods can be broadly categorized into two groups: metapath-based and metapath-free approaches. Metapath-based methods fu2020magnn ; zhang2019heterogeneous ; wang2019heterogeneous ; schlichtkrull2018modeling ; dong2017metapath2vec ; sun2011pathsim ; yun2019graph ; yang2023simple ; zhu2024hagnn first aggregate features from neighbors sharing the same semantic context and then integrate information across different semantics. On the other hand, metapath-free methods hu2020heterogeneous ; lv2021we ; zhu2019relation ; hong2020attention ; zhang2022simple ; fu2023multiplex ; huo2025heterogeneous aggregate information from all types of neighbors within a local 1-hop neighborhood, akin to traditional GNNs, but incorporate additional mechanisms such as attention to encode semantic details (e.g., node and edge types) into the message-passing process. While HGNNs have achieved remarkable success in learning from static heterogeneous graphs, they typically rely on the assumption that the entire graph is accessible during training. This assumption limits their applicability in dynamic real-world scenarios, where graph structures evolve continuously over time, as seen in knowledge graphs and social networks. Consequently, Continual Learning (CL) emerges as a critical yet underexplored research direction for advancing HGNNs in such dynamic settings.
@@
Continual Graph Learning. Continual Graph Learning (CGL) enables models to learn from evolving graph-structured data while mitigating catastrophic forgetting, where performance on earlier tasks deteriorates after learning new ones. Existing methods can be broadly categorized into three types: parameter-isolation, regularization, and memory-replay approaches. Parameter-isolation methods preserve important parameters by freezing them after training on a task, preventing interference from subsequent updates. Examples include HPNs zhang2022hierarchical and the PI-GNN framework zhang2023continual . Regularization methods constrain updates to important parameters via penalty terms. Representative methods include EWC kirkpatrick2017overcoming , MAS Aljundi_2018_ECCV , and TWP liu2021overcoming , which considers local graph topology. Knowledge distillation is also commonly used to retain information from prior models Xu2020GraphSAIL ; dong2021rkd . Memory-replay methods sample and store representative nodes or subgraphs for replay zhang2023ricci ; ahrabian2021structure . However, these methods may suffer from memory explosion due to the growth of computational subgraphs. ER-GNN zhou2021overcoming addresses this by sampling node attributes only, while SSM zhang2022sparsified stores sparsified subgraphs. PDGNNs-TEM zhang2024topology maintains dynamic embeddings of subgraphs, and UGCL hoang2023universal unifies memory replay and distillation to support both node and graph classification tasks. TACO han2024topology replaces raw subgraph storage with graph coarsening, enabling compact memory usage while preserving key structural information. A recent work, RL-GNN song2025exploring , addresses graph-level continual learning by identifying invariant rationales across tasks to mitigate catastrophic forgetting caused by spurious correlations. Despite recent progress, CGL still faces challenges in reducing forgetting and computational cost. Moreover, continual learning on heterogeneous graphs remains underexplored. In this work, we address these challenges using a combination of meta-learning and knowledge distillation tailored to heterogeneous graph settings.
@@
Meta Learning. Meta-learning aims to enable models to adapt rapidly to new tasks by leveraging shared knowledge. Existing approaches are generally categorized into optimization-based (e.g., MAML finn2017model , Reptile nichol2018reptile ) and metric-based methods (e.g., Prototypical Networks snell2017prototypical , Matching Networks vinyals2016matching ). In continual learning, meta-learning has been applied to online gupta2020look and few-shot scenarios javed2019meta . Recently, its integration with graph learning has drawn increasing attention. For example, MetaCLGraph unal2023meta incorporates meta-learning with experience replay for continual graph learning, and HG-Meta zhang2022hg applies it to few-shot graph classification on heterogeneous graphs. Motivated by these advances, we further explore its potential in mitigating forgetting and improving adaptability in heterogeneous continual learning.
3 Preliminary
@@
In this section, we introduce fundamental concepts and formalize the problem of Heterogeneous Continual Graph Learning (HCGL). Our primary objective is to develop a continual learning framework on Dynamic Heterogeneous Graphs to mitigate the issue of catastrophic forgetting. In the HCGL setting, a Heterogeneous Graph Neural Network (HGNN) learns a sequence of tasks in a domain incremental setting zhang2024continual (See Appendix B.3 for the details of this setting), without access to the data from previous tasks. However, it is allowed to utilize a memory buffer with limited capacity to store representative information. The goal of the framework is to maximize the prediction accuracy across all tasks after training while minimizing the forgetting of previously acquired knowledge during the learning of new tasks.
@@
Let $\mathcal{G}=(V,E)$ be a dynamic heterogeneous graph, where $V$ is the node set and $E$ is the edge set. The feature set is type-specific, i.e., =$\mathcal{F}=\{F_{\tau}\in\mathbb{R}^{|V_{\tau}|\times d_{\tau}}\mid\tau\in\mathcal{A}\}$. In the HCGL setting, the model learns a sequence of tasks $\{\mathcal{T}_{1},\mathcal{T}_{2},\dots,\mathcal{T}_{T}\}$, without access to the data from previous tasks. Specifically, each task $\mathcal{T}_{t}$ corresponds to a subgraph $\mathcal{G}_{t}$ with disjoint category labels, and due to storage limitations, the model can only access the data available at the current task. For each subgraph $\mathcal{G}_{t}$, we divide it into a training set $\mathcal{G}_{t}^{\text{tr}}=(V_{t}^{\text{tr}},E_{t}^{\text{tr}})$ and a testing set $\mathcal{G}t^{\text{te}}=(V_{t}^{\text{te}},E_{t}^{\text{te}})$ to train and evaluate the model $f_{\theta}$.
4 Methodology
In this section, we present our Meta-learning based Knowledge Distillation (MKD) framework, providing a detailed analysis of its core components and demonstrating how they systematically address the identified challenges.
Figure 1: The overall framework of MKD. (a) Subgraph Sampling: Construct task-specific subgraphs by selecting diverse target-type nodes and expanding to related node types via relation-aware importance. (b) Knowledge Distillation: Align previous and current tasks via logit-level and semantic-level distillation. (c) Task Training: Use meta-learning for fast adaptation, jointly optimizing task loss, replay loss, and distillation loss.
4.1 Fast Adaptation with Gradient-based Meta-Learning
Efficient adaptation to new data patterns is crucial for ensuring the practical usability of a model in an evolving heterogeneous graph. To achieve this, we introduce the Gradient-based Meta-learning Module (G-MM), which optimizes model parameter initialization to enable rapid adaptation to new tasks. This, in turn, helps preserve performance on the current task. Specifically, given the training node set $V_{t}^{tr}$ of the current task $\mathcal{T}_{t}$, we select $e$ samples from $V_{t}^{tr}$ based on the Coverage Maximization (CM) strategy (introduced in Section 4.2), which identifies key nodes by maximizing node diversity. In practice, labeled data in heterogeneous graphs is often highly limited. By performing gradient descent on the selected small sample set $\mathcal{E}$, the model can quickly adapt to the current task, a process referred to as the inner update. Let $\theta$ denote the set of model parameters. For the current task $\mathcal{T}_{i}$, We feed the sampled node set $\mathcal{E}$ into the model and compute the loss $\mathcal{L}_{\mathcal{E}}$, updating $\theta$ to $\theta_{fast}$ via gradient descent (Inner Update):
$\theta_{fast}\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\mathcal{E}}(\theta),$ (1)

### v2

# HERO: Heterogeneous Continual Graph Learning via Meta-Knowledge Distillation

@@
Abstract.
Heterogeneous graph neural networks have seen rapid progress in web applications such as social networks, knowledge graphs, and recommendation systems, driven by the inherent heterogeneity of web data. However, existing methods typically assume static graphs, while real-world graphs are continuously evolving. This dynamic nature requires models to adapt to new data while preserving existing knowledge. To this end, this work introduces HERO (HEterogeneous continual gRaph learning via meta-knOwledge distillation), a unified framework for continual learning on heterogeneous graphs. HERO employs meta-adaptation, a gradient-based meta-learning strategy that provides directional guidance for rapid adaptation to new tasks with limited samples. To enable efficient and effective knowledge reuse, we propose DiSCo (Diversity Sampling with semantic Consistency), a heterogeneity-aware sampling method that maximizes target node diversity and expands subgraphs along metapaths, retaining critical semantic and structural information with minimal overhead. Furthermore, HERO incorporates heterogeneity-aware knowledge distillation, which aligns knowledge at both the node and semantic levels to balance adaptation and retention across tasks. Extensive experiments on four web-related heterogeneous graph benchmarks demonstrate that HERO substantially mitigates catastrophic forgetting while achieving efficient and consistent knowledge reuse in dynamic web environments.
Incremental Learning, Model Reuse, Stability-Plasticity Trade-off
Corresponding authors: Xikun Zhang and Dongjin Song
$c$s: Computing methodologies Lifelong machine learningccs: Information systems Social networks
1. Introduction
@@
The modern web is inherently heterogeneous, consisting of diverse entities and relations such as users, items, hyperlinks, and knowledge facts. As a result, heterogeneous graphs (HGs) have become a natural abstraction for modeling complex web systems including social networks, knowledge graphs, and recommendation platforms (Shi et al., 2018; Sun and Han, 2012). Unlike homogeneous graphs with uniform node and edge types, HGs encode multifaceted relationships across different types of nodes and edges, enabling rich and expressive representations for web-scale applications. For example, in recommendation systems, user–item interactions form user nodes, item nodes, and purchase edges; in knowledge graphs, entities and relations compose semantic networks that support web search and reasoning; in social media, dynamic interactions between users, posts, and topics constitute evolving heterogeneous graphs that drive personalized content delivery. To extract meaningful insights from heterogeneous graphs, different Heterogeneous Graph Neural Networks (HGNNs) have been developed. These approaches mainly fall into two categories: metapath-based HGNNs (Fu et al., 2020; Zhang et al., 2019; Wang et al., 2019; Schlichtkrull et al., 2018; Dong et al., 2017; Sun et al., 2011; Yun et al., 2019; Yang et al., 2023; Zhu et al., 2024), which aggregate information over predefined meta-paths to mine the intricate relationship among the heterogeneous nodes and edges, and meta-path-free models (Hu et al., 2020; Lv et al., 2021; Zhu et al., 2019; Hong et al., 2020; Zhang et al., 2022e; Fu et al., 2023; Huo et al., 2025) that automatically learn node interactions without preconfigured paths, offering enhanced adaptability but potentially sacrificing interpretability.
Figure 1. An example of Evolving Heterogeneous Graph. (a) Node types include Actor, Movie, and Director. (b) As the graph evolves, new domains emerge (e.g., movies of different genres such as Comedy/Romance and Thriller/Action).
Despite the success of Heterogeneous Graph Neural Networks (HGNNs) in static settings, real-world web graphs are inherently evolving. In movie-related graphs such as IMDB, new entities and relations emerge over time, reflecting shifts in genres, collaborations, and audience interests (Figure 1). In such evolving environments, HGNNs are expected to continuously incorporate the incoming new knowledge while preserving the learnt patterns. However, existing continual learning (CL) methods, largely developed for images (Aljundi et al., 2018; Kirkpatrick et al., 2017; Li and Hoiem, 2017; Lopez-Paz and Ranzato, 2017) and homogeneous graphs (Zhang et al., 2022a; Liu et al., 2021; Zhou and Cao, 2021; Zhang et al., 2023b; Unal et al., 2023; Su et al., 2023; Ren et al., 2023; Niu et al., 2024; Li et al., 2024; Mondal et al., 2024; Liu et al., 2024; Choi et al., 2024; Qiao et al., 2025), fall short when directly applied to heterogeneous graphs due to semantic diversity and structural imbalance. For instance, in recommendation platforms, rare product categories are easily forgotten; in knowledge graphs, semantically distinct relations collapse into uniform representations; and in social networks, global regularization fails to capture type-specific parameter importance. These challenges suggest the necessity to develop new continual learning strategies which can explicitly account for web-scale heterogeneity.
@@
To bridge this critical gap, we present a systematic investigation of the Heterogeneous Continual Graph Learning (HCGL) problem and introduce HERO (HEterogeneous continual gRaph learning via meta-knOwledge distillation), a unified framework for continual learning on heterogeneous graphs. Unlike homogeneous graphs, the diverse node and edge types in heterogeneous graph introduce extra challenges to knowledge transfer across different tasks. Therefore, for effective model adaptation over the continuously evolving heterogeneous graph structures, we design a Gradient-based Meta-learning Module (G-MM). G-MM learns an optimized initialization of model parameters by leveraging gradient-based meta-learning, which allows fast model adaptation to new tasks using only a small number of examples. Through task-specific adjustments, G-MM also avoids the task-wise interference and ensures performance robustness on learned tasks. Memory replay has proven to be an effective strategy for mitigating catastrophic forgetting. However, existing replay methods typically rely on storing large volumes of historical data, which makes continual learning impractical in large-scale tasks. Moreover, conventional memory sampling strategies cannot be directly applied to HCGL due to the complexity of heterogeneous structures. To address this, we propose DiSCo (Diversity Sampling with semantic Consistency), a heterogeneity-aware sampling method that maximizes the diversity of target nodes and expands subgraphs along meta-paths, thereby retaining critical semantic and structural information with minimal memory overhead. Furthermore, to align knowledge across tasks and maintain semantic consistency in heterogeneous settings, we introduce a Heterogeneity-aware Knowledge Distillation (HKD) module. Beyond knowledge retained via experience replay, this module performs both logit-level and semantic-level alignment between the current and previous task models. By distilling prediction distributions and meta-path-based attention representations, HKD guides the student model to retain predictive information while incorporating multi-level semantic information, significantly improving continual learning performance on heterogeneous graphs. Our contributions are summarized as follows:
  1. (1)
    We provide the first systematic investigation of the Heterogeneous Continual Graph Learning (HCGL) problem, highlighting its unique challenges in web-scale settings compared to homogeneous graphs.
  2. (2)
@@
    We propose DiSCo, a diversity- and semantics-aware sampling strategy that enables memory-efficient and structurally balanced experience replay.
  3. (3)
@@
    We design a heterogeneity-aware knowledge distillation mechanism that aligns knowledge at both logit and semantic levels to balance adaptation and retention.
  4. (4)
    We develop the HERO framework, a novel framework tailored for HCGL. Experiments on multiple heterogeneous web graph benchmarks (e.g., knowledge graph, recommendation and citation networks datasets) validate its superior performance in terms of accuracy, efficiency, and robustness.
2. Related Work
Heterogeneous Graph Neural Networks.
@@ 2. Related Work
Heterogeneous graphs are prevalent in web environments such as knowledge graphs, social networks, and recommendation systems, where multiple node and edge types naturally coexist. These graphs are characterized by complex topological structures and rich semantic information, which traditional Graph Neural Networks (GNNs) struggle to model directly. To address this challenge, specialized approaches known as Heterogeneous Graph Neural Networks (HGNNs) have been developed to effectively capture the multi-typed relationships inherent in web graphs. Existing HGNN methods can be broadly categorized into two groups: metapath-based and metapath-free approaches. Metapath-based methods (Fu et al., 2020; Zhang et al., 2019; Wang et al., 2019; Schlichtkrull et al., 2018; Dong et al., 2017; Sun et al., 2011; Yun et al., 2019; Yang et al., 2023; Zhu et al., 2024) explicitly leverage predefined semantic paths, aggregating features from neighbors within the same semantic context and integrating across metapaths. In contrast, metapath-free methods (Hu et al., 2020; Lv et al., 2021; Zhu et al., 2019; Hong et al., 2020; Zhang et al., 2022e; Fu et al., 2023; Huo et al., 2025) adopt a more flexible design by aggregating information from all types of neighbors within a local neighborhood, while encoding semantic distinctions (e.g., node and edge types) through mechanisms such as attention. These methods have achieved remarkable success on static web graphs; however, they typically assume full access to the entire graph during training. This assumption becomes restrictive in realistic web scenarios, where graphs evolve continuously with new nodes and relations. To this end, Continual Learning (CL) for HGNNs emerges as a crucial yet underexplored research direction, enabling models to adapt to dynamic web graphs while preserving previously acquired knowledge.
Continual Graph Learning.
@@
Continual Graph Learning (CGL) enables models to learn from evolving graph-structured data while mitigating catastrophic forgetting, where performance on earlier tasks deteriorates after learning new ones. Existing methods can be broadly categorized into three types: parameter-isolation, regularization, and memory-replay approaches. Parameter-isolation methods preserve important parameters by freezing them after training on a task, preventing interference from subsequent updates. Examples include HPNs (Zhang et al., 2022b) and the PI-GNN framework (Zhang et al., 2023b). Regularization methods constrain updates to important parameters via penalty terms. Representative methods include EWC (Kirkpatrick et al., 2017), MAS (Aljundi et al., 2018), and TWP (Liu et al., 2021), which considers local graph topology. Knowledge distillation is also commonly used to retain information from prior models (Xu et al., 2020; Dong et al., 2021). Memory-replay methods sample and store representative nodes or subgraphs for replay (Zhang et al., 2023a; Ahrabian et al., 2021). However, these methods may suffer from memory explosion due to the growth of computational subgraphs. ER-GNN (Zhou and Cao, 2021) addresses this by sampling node attributes only, while SSM (Zhang et al., 2022c) stores sparsified subgraphs. DSLR (Choi et al., 2024) further enhances replay by promoting sample diversity and learning informative subgraph structures to better preserve past knowledge. PDGNNs-TEM (Zhang et al., 2024b) maintains dynamic embeddings of subgraphs, and UGCL (Hoang et al., 2023) unifies memory replay and distillation to support both node and graph classification tasks. TACO (Han et al., 2024) replaces raw subgraph storage with graph coarsening, enabling compact memory usage while preserving key structural information. A recent work, RL-GNN (Song et al., 2025), addresses graph-level continual learning by identifying invariant rationales across tasks to mitigate catastrophic forgetting caused by spurious correlations. Despite recent progress, CGL still faces challenges in reducing forgetting and computational cost. Moreover, continual learning on heterogeneous graphs remains underexplored. In this work, we address these challenges by using proposed DiSCo and Heterogeneity-aware Knowledge Distillation tailored to heterogeneous graph settings.
Meta Learning.
@@
Meta-learning aims to enable models to adapt rapidly to new tasks by leveraging shared knowledge. Existing approaches are generally categorized into optimization-based (e.g., MAML (Finn et al., 2017), Reptile (Nichol and Schulman, 2018)) and metric-based methods (e.g., Prototypical Networks (Snell et al., 2017), Matching Networks (Vinyals et al., 2016)). By learning how to learn, meta-learning provides a principled way to capture transferable inductive biases that facilitate fast adaptation to new tasks with limited data. In the context of continual learning, meta-learning has been applied to both online (Gupta et al., 2020) and few-shot settings (Javed and White, 2019), where it helps alleviate catastrophic forgetting by enabling rapid re-optimization when task distributions shift. Recently, the integration of meta-learning with graph neural networks has drawn increasing attention, as graphs naturally exhibit diverse and evolving structures that demand adaptive learning capabilities. For example, MetaCLGraph (Unal et al., 2023) incorporates meta-learning with experience replay to balance adaptation and stability across sequential graph tasks, while HG-Meta (Zhang et al., 2022d) extends meta-learning to few-shot scenarios on heterogeneous graphs by modeling task-level semantics and relational dependencies. Motivated by these advances, we further explore the potential of meta-learning in mitigating forgetting and improving adaptability under heterogeneous continual graph learning, where both structural and semantic variations pose additional challenges compared to conventional settings.
3. Preliminary
@@
In this section, we introduce fundamental concepts and formalize the problem of Heterogeneous Continual Graph Learning (HCGL). Our primary objective is to develop a continual learning framework on Evolving Heterogeneous Graphs to mitigate the issue of catastrophic forgetting. In the HCGL setting, a Heterogeneous Graph Neural Network (HGNN) learns a sequence of tasks in a domain incremental setting (Zhang et al., 2024a) (See Appendix B.3 for the details of this setting), without access to the data from previous tasks. However, it is allowed to utilize a memory buffer with limited capacity to store representative information. The goal of the framework is to maximize the prediction accuracy across all tasks after training while minimizing the forgetting of previously acquired knowledge during the learning of new tasks.
@@
Let $\mathcal{G}=(V,E)$ be a dynamic heterogeneous graph, where $V$ is the node set and $E$ is the edge set. The feature set is type-specific, i.e., $\mathcal{F}=\{F_{\tau}\in\mathbb{R}^{|V_{\tau}|\times d_{\tau}}\mid\tau\in\mathcal{A}\}$. In the HCGL setting, the model learns a sequence of tasks $\{\mathcal{T}_{1},\mathcal{T}_{2},\dots,\mathcal{T}_{T}\}$, without access to the data from previous tasks. Specifically, each task $\mathcal{T}_{t}$ corresponds to a subgraph $\mathcal{G}_{t}$ with disjoint category labels, and due to storage limitations, the model can only access the data available at the current task. For each subgraph $\mathcal{G}_{t}$, we divide it into a training set $\mathcal{G}_{t}^{\text{tr}}=(V_{t}^{\text{tr}},E_{t}^{\text{tr}})$ and a testing set $\mathcal{G}t^{\text{te}}=(V_{t}^{\text{te}},E_{t}^{\text{te}})$ to train and evaluate the model $f_{\theta}$.
4. Methodology
In this section, we present our HERO (HEterogeneous continual gRaph learning via meta-knOwledge distillation) framework (see Figure 2), provide a detailed introduction of its core components, and explain how it can systematically address the identified challenges.
Figure 2. The overall framework of HERO. (a) Diversity Sampling with semantic Consistency: Construct task-specific subgraphs by selecting diverse target-type nodes and expanding to related node types via relation-aware importance. (b) Heterogeneity-aware Knowledge Distillation: Align previous and current tasks via logit-level and semantic-level distillation. (c) Task Training: Use meta-learning for fast adaptation, jointly optimizing task loss, replay loss, and distillation loss.
4.1. Fast Adaptation with Gradient-based Meta-Learning
Efficient adaptation to new data patterns is crucial for ensuring the practical usability of a model in an evolving heterogeneous web graph. To achieve this, we introduce the Gradient-based Meta-learning Module (G-MM), which optimizes model parameter initialization to enable rapid adaptation to new tasks. This, in turn, helps preserve performance on the current task. Specifically, given the training node set $V_{t}^{tr}$ of the current task $\mathcal{T}_{t}$, we select $e$ samples from $V_{t}^{tr}$ based on the Coverage Maximization (CM) strategy (introduced in Section 4.2), which identifies key nodes by maximizing node diversity. In practice, labeled data in heterogeneous graphs is often highly limited. By performing gradient descent on the selected small sample set $\mathcal{E}$, the model can quickly adapt to the current task, a process referred to as the inner update. Let $\theta$ denote the set of model parameters. For the current task $\mathcal{T}_{i}$, We feed the sampled node set $\mathcal{E}$ into the model and compute the loss $\mathcal{L}_{\mathcal{E}}$, updating $\theta$ to $\theta_{fast}$ via gradient descent (Inner Update):
(1) $\theta_{fast}\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\mathcal{E}}(\theta),$